OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4 A.) -4 and 0 B.)4 and 0 C.) -2 and -3 D.) 2 and -3

OpenStudy (anonymous):

@wio

myininaya (myininaya):

Did you try replacing y in the first equation with x+4?

OpenStudy (anonymous):

Im not sure if I did it correctly but I got B @myininaya

myininaya (myininaya):

If you think it is B, you can check it by pluging your answer into the equations given. If both equations are satisfied then you did good.

myininaya (myininaya):

Well first you would need to find y though for both x's

myininaya (myininaya):

You can use that bottom equation. It is the most easiest.

OpenStudy (anonymous):

You don't need to actually solve the equation. In this question you can just substitute in the answers for \(x\) and see which answer is consistent.

myininaya (myininaya):

So you have x=4 or x=0 If x=4, then y=4+4=8 (Plug in (4,8)) If x=0, then y=0+4=4 (Plug in (0,4))

OpenStudy (anonymous):

I re-did it and got A as my answer

myininaya (myininaya):

Cool stuff. That answer is actually correct.

OpenStudy (anonymous):

thank you guys @myininaya @wio !